OpenStudy (anonymous):

F(x) = (1/x-1) , x>1 , a=2 a- Show that f is one-to-one. b- Find (f^-1)’(a). c- Calculate (f^-1)(x) and state the domain and range of (f^-1). d- Calculate (f^-1)’(a) from the formula in part (c) and check that it agrees with the result of part (b). e- Sketch the graphs of f and (f^-1) on the same axes. plzz help me !

OpenStudy (anonymous):

a) f(x1)=f(x2) => x1=x2 is basically the definition for one-to-one function (mind you both x1 and x2 should be in domain)... 1/(x1-1) = 1/(x2-1) => x1=x2... hence, it is one-to-one...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b) as our function's Range is (0,inf) we see that it is onto as well... so, inverse exists... f(x)=y => f^-1(y) =x -----> (1) 1/(x-1) = y => x= (1/y)+1 ----> (2) from (1) and (2) => f^-1(y) = (1/y)+1... f^-1(2) = 1.5 ...

OpenStudy (anonymous):

srry, my mistake... thts ur answer for part (c)

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

i dont understand the part (b) but everything else can be found from answer to part (c) ... is that enough ??

OpenStudy (anonymous):

amm what about part d its like part c ?

OpenStudy (anonymous):

is it*

OpenStudy (anonymous):

now that u have got the function find its derivative... using standard rules

OpenStudy (anonymous):

aha ok

OpenStudy (anonymous):

thanks